OpenStudy (anonymous):

What is the graph of the function f(x) =-x^2+x+20/x+4

OpenStudy (anonymous):

Start by factoring the top and bottom.

OpenStudy (anonymous):

how do i do that?

OpenStudy (solomonzelman):

it is a line. Factor the top.

OpenStudy (anonymous):

Let me give you an example : x^2 +6x +4 This factors to (x+4)(x+2)

OpenStudy (solomonzelman):

x+4 cancel on top and bottom

OpenStudy (jdoe0001):

\(\bf -x^2+x+20\implies -(x^2-x-20)\)

OpenStudy (jdoe0001):

thus factor the numerator.... and \(\bf f(x)=\cfrac{(\qquad )\cancel{ (\qquad ) }}{\cancel{ x+4 }}\)

OpenStudy (anonymous):

-(+4)(x-5)?

OpenStudy (jdoe0001):

-[ (x+4)(x-5) ] is correct so \(\bf f(x)=\cfrac{-[\cancel{ ( x+4) }(x-5 )]}{\cancel{ x+4 }}\)

OpenStudy (jdoe0001):

so that'd be the graph of that is a line, pick a couple of random "x" and get a "y" plot the line :) keeping in mind the "restrictions" of the original rational that when the denominator is 0, the fraction becomes "undefined" and thus has a hole there due to discontinuity \(\bf f(x)=\cfrac{-x^2+x+20}{{\color{brown}{ x+4}}}\qquad {\color{brown}{ x+4}}=0\qquad \cfrac{-x^2+x+20}{{\color{brown}{ 0}}}\to discontinuous\)

OpenStudy (jdoe0001):

hmm actually..... it may not be discontinuous

OpenStudy (solomonzelman):

\(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid white ]{ \LARGE\text{Chart    for    f(x) .}\\ \begin{array}{|c|c|c|c|} \hline~~~~~~\textbf{x}~~~~~~ ~&~~~~~~\textbf{f(x) }~~~~~~\\ \hline\text{0} &\text{4} \\ \text{1} &\text{5} \\ \text{2} &\text{6} \\ \text{3} &\text{7} \\ \text{4} &\text{8} \\ \text{••••} &\text{••••} \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

well, jdoe, it is just f(x)=x+4 , technically.

OpenStudy (jdoe0001):

yeah.... when the denominator zeros out, so does the numerator....so it wouldn't be discontinous

OpenStudy (anonymous):

none of the graphs that are answers are just like a straight line... im confused